# SuperGSeg: Open-Vocabulary 3D Segmentation with Structured Super-Gaussians

### v2

@@ 3.3 Super-Gaussian Clustering
$\widehat{\mathbf{e}}_{j}^{t+1}=\frac{1}{\sum_{j=i}^{N^{\prime}}\mathbb{I}(j\in\mathcal{N}_{i}){{\bm{A}}}_{ij}^{t}}\sum_{i=1}^{N^{\prime}}\mathbb{I}(j\in\mathcal{N}_{i}){{\bm{A}}}_{ij}^{t}\mathbf{e}_{i},$ (4)
@@
We train the association probability matrix $\bm{A}$}} and the SuperG s’ attributes $\widehat{\mathbf{e}}$ with anchors’ attributes $\mathbf{e}$, where anchors attributes $\mathbf{e}$ are optimized during Stage 1 (Section 3.2) and frozen in this stage:
@@
$\mathcal{L}_{compact,\mathcal{X}}=\frac{1}{S}\sum_{j=1}^{S}\lVert\mathcal{X}_{j}-\widehat{{\bm{x}}}_{j}\rVert,$ (6)
@@ 4.1 Experimental Setup
Baselines and Metrics. We compare our method with representative NeRF-based and 3DGS-based baselines, including LERF [13], LangSplat [29], LEGaussian [34], and OpenGaussian [38]. For the open-vocabulary semantic segmentation task, CLIP-encoded text features are compared with rendered 2D language feature maps via cosine similarity to produce per-pixel semantic predictions [43], evaluated with mean Intersection over Union (mIoU) and mean Accuracy (mAcc). For the open-vocabulary object selection task, we perform text queries directly in 3D space [38], retrieving the most relevant SuperG and rendering them into 2D for evaluation with mIoU and mAcc. Since NeRF is an implicit representation without explicit 3D positions, LERF cannot be applied to this task. We also report inference-time efficiency, measuring both runtime and memory consumption for text queries on trained 3D scenes. Specifically, we perform multiple queries from different viewpoints and report the average query time. We consider this metric particularly important for assessing the feasibility of deploying models on resource-constrained devices and enabling real-time querying in practical scenarios.
@@ 6 Acknowledgement
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP 02, project number: 276693517. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Siyun Liang.
@@
This supplementary document provides additional details about our method. Section A elaborates on the design of the Super-Gaussian and demonstrates its application to downstream tasks. Section B provides further implementation details, including the MLP architectures for neural Gaussian feature decoding and the adaptation of OpenGaussian for 2D open-vocabulary semantic segmentation comparison. Section D reports extended ablation studies on SuperG hyperparameters and module variants. Section E presents additional quantitative and qualitative results. Lastly, Section F discusses the limitations of our approach and outlines potential directions for future work.
@@ Appendix A Super-Gaussian Details
Figure 8: Examples of Super-Gaussian graph. Each node represents a SuperG, connected to its $k$-nearest neighbors based on similarity in instance features. Through connected component analysis, the SuperG nodes are divided into two distinct instances. Within Instance 1, SuperG nodes are interconnected by similar hierarchical features, further splitting into two parts.
Module Design. As shown in Figure 3, our SuperG clustering network consists of four learnable MLPs. Inspired by SPNet [10], we design three attribute-specific learnable functions $F_{\phi}$, $F_{\varphi}$, and $F_{\psi}$, each implemented as a single-hidden-layer MLP with ReLU activation. These functions independently encode the differences between an anchor and its $k$-nearest SuperG s in coordinates, segmentation features, and geometry, respectively, producing embeddings that reflect the relevance of each attribute for SuperG assignment. A final MLP, $F_{sg}$, then concatenates these embeddings and integrates spatial, semantic, and geometric cues into probabilistic assignment.
@@
After training the SuperG association modules, we obtain a soft association map ${\bm{A}}\in\mathbb{R}^{N^{\prime}\times k}$. During inference, each anchor point is assigned to one of its $k$-nearest neighbors with the highest probability, leading to a hard SuperG assignment $\bar{{\bm{A}}}\in\mathbb{R}$. The attributes of each SuperG are then computed by averaging the attributes of its assigned anchors.
These SuperG s serve as the fundamental units for representing and interpreting the 3D scene. Specifically, as shown in Figure 8, we further construct a graph where nodes correspond to SuperG s. For instance segmentation, a node is connected to nodes within its $k$ nearest neighbors if their instance feature similarity exceeds a threshold $\tau_{Ins}$. Instances are then obtained via connected component analysis on this restricted graph. Similarly, part segmentation is achieved by building a SuperG graph within each instance and identifying connected components based on hierarchical feature similarity with threshold $\tau_{Hier}$. In practice, we set $k=3$, $\tau_{Ins}=0.8$, and $\tau_{Hier}=0.9$.
@@ Appendix B Additional Implementation Details
Figure 9: MLP structures to decode different features.
@@
We employ MLPs to decode latent features, as shown in Figure 9. Each MLP contains a single hidden layer of dimension 32. Their decoding targets, however, differ: $F_{I}$ and $F_{H}$ decode anchor-level features, while $F_{L}$ decodes SuperG-level features. Specifically, $F_{I}$ and $F_{H}$ take the anchor segmentation feature ${\bm{f}}^{s}$ and anchor position $\bm{x}$} as input, and predict the instance feature $\bm{g}$} and hierarchical feature $\bm{h}$} of the neural Gaussians spawned per anchor. In contrast, $F_{L}$ predicts the CLIP-aligned feature $\widehat{{\bm{f}}}^{l}$ for each SuperG, conditioned on its latent language feature $\widehat{{\bm{f}}}^{l}$ and its center $\widehat{{\bm{x}}}$.
@@
The contrastive instance feature loss $\mathcal{L}_{Ins}$, which is computed on the set of SAM-generated instance-level masks $\mathcal{M}$ and the rendered 2D instance feature map $\hat{{\bm{G}}}$, has been discussed. Inspired by [41], we elaborate a similar yet more complicated contrastive hierarchical feature loss $\mathcal{L}_{Hier}$. This loss is defined on the part-level patches $\mathcal{P}=\{{{\bm{p}}}^{p}\in{\mathbb{R}}^{H\times W}\mid p=1,\dots,|\mathcal{P}|\}$ and the rendered 2D hierarchical feature map $\hat{{\bm{H}}}\in{\mathbb{R}}^{D_{h}\times H\times W}$.
@@
We use the SAM ViT-H model to generate 2D masks from the input images and then extract language features for each instance mask using the OpenCLIP ViT-B/16 model following [29]. The training process is divided into three stages. In the first stage, we train the Scaffold-GS [23] with instance and hierarchical feature attributes for 30k iterations. In the second stage, we freeze the geometry and multi-granularity features network from stage one and train only the SuperG clustering network for another 30k iterations. Finally, in the last stage, we freeze all other parameters and optimize the language features for each SuperG for 10k iterations. The embedding dimensions for ${\bm{f}}^{g}$ and ${\bm{f}}^{s}$ are set to 32 [23], while instance and hierarchical features are 16-dimensional [41]. For optimization, we use Adam [15] optimizer for the MLPs with an initial learning rate of 0.01 and an exponential annealing schedule of 0.001 as in [8].
@@
OpenGaussian [38] assigns language features to instance-level Gaussian, enabling direct language queries on 3D point clouds. However, this approach does not natively support 2D pixel-level semantic segmentation, making direct evaluation on ScanNet more challenging. To enable a fair comparison, we first identify category-relevant 3D Gaussians by iterating over all text prompts to predict language feature maps for open-vocabulary semantic segmentation. For each instance-level Gaussian cluster, we determine the corresponding text prompt ID and store these IDs in a label map, which is then used to generate the final semantic segmentation. By following this approach, occlusions at the instance-level Gaussian are not explicitly handled, leading to the occlusion artifacts observed in Figure 4.
@@
Table 4: Additional ablation studies on the LERF-OVS dataset [29] about the parameter choice for SuperG Clustering Network. We use $s=1000$ SuperG s and $k=3$ fr $k$-nearest neighbor in our implementation by default.
Appendix C Training Efficiency and Inference Advancement
@@
In Table 5, we report training time, inference time and memory consumption with LangSplat [29] and OpenGaussian [38] on the LERF-OVS dataset.
@@
In comparing our method to LangSplat and OpenGaussian across different training stages S1, S2, and S3, we find that our approach, while requiring longer training times in S1 and S2, achieves comparable efficiency in S3, indicating a trade-off due to the joint learning of instance and hierarchical features. GPU memory usage varies, with our method consuming more resources in S1 and S2 but significantly less in S3, showcasing the compact memory footprint enabled by our proposed SuperG s. Conversely, LangSplat exhibits consistent memory efficiency in S1 and S2, while OpenGaussian’s memory demands vary across different training stages. Most notably, our method consistently outperforms both baselines in inference speed across all scenarios, a critical advantage for real-time applications. This blend of rapid inference and flexible resource utilization highlights our method’s robustness for practical deployment in diverse computer vision tasks. In addition, our method supports multi-granularity scene understanding. Specifically, it enables semantic-level queries to retrieve groups of objects sharing the same language description, instance-level segmentation of a specific object, and further decomposition of this instance into fine-grained parts. In contrast, LangSplat [29] only supports semantic-level queries. OpenGaussian [38] extends to instance-level understanding but, similar to LangSplat, does not support finer-grained part segmentation. This demonstrates that our method provides a more comprehensive representation for 3D scene understanding.
@@ Appendix D Additional Ablation Studies
Super-Gaussian Generation Approaches.
Given a pre-trained scene, our objective is to group anchor into meaningful SuperG s using their coordinates, segmentation features, and geometric properties. We explore two alternative approaches. First, we evaluate a simple $K$-Means clustering algorithm by concatenating the aforementioned attributes and clustering them into $k=1000$ SuperG s. Second, we experiment with a traditional supervoxel generation approach. Specifically, we map the anchors to a point cloud, using the concatenated features as normals. We then apply the Voxel Cloud Connectivity Segmentation (VCCS) algorithm [26] to compute the SuperG s. Finally, we compare these two non-learning-based approaches with our learning-based method for Anchor-to-SuperG association.
@@
Table 6: SuperG clustering approaches ablation study on the teatime scene of LERF-OVS.
@@
Table 7: SuperG Clustering Network and components ablation study on the teatime scene of LERF-OVS.
@@
As introduced in Section A, we employ three MLPs $F_{\phi}$, $F_{\varphi}$ and $F_{\psi}$, to capture the coordinate, segmentation, and geometric relationships between anchors and their $k$ nearest neighbors. To evaluate the contributions of these MLPs, we conduct an ablation study. Notably, in experiments where none of these MLPs are used, we directly concatenate the differences of the attributes as input to $F_{sg}$ for predicting the association matrix. The results presented in Table 7 demonstrate that each MLP contributes to improving the SuperG assignments. The MLP $F_{\varphi}$, which accounts for the segmentation feature differences between the anchor and the SuperG, has the most significant impact. Notably, using only $F_{\varphi}$ yields a relatively high mIoU, emphasizing its effectiveness in aligning semantic features. However, our full setup that integrates the $F_{\phi}$ and $F_{\psi}$ for coordinate and geometric feature information further enhances mAcc. This suggests that incorporating additional spatial and geometric context refines the SuperG assignments, leading to more precise understanding of the scene.
@@ Appendix E Additional Results
Qualitative Result in Occlusion Cases.
As illustrated in Figure 10, our method queries objects directly from 3D space, effectively mitigating occlusion issues (the bear leg under the table can be retrieved). Moreover, the queried objects exhibit multi-view consistency, enabling comprehensive scene understanding in 3D. For additional qualitative results, please refer to the accompanying videos.
Additional Quantitative Result on ScanNet.
We report the results on more categories in ScanNet dataset in Table 8.
@@ Appendix F Limitations
Despite the advancements achieved by our method, certain limitations remain. First, our approach inherits biases from the original visual foundation models, which constrains performance to a local maximum and may limit the generalization of results to diverse or unseen scenarios. Second, our method is tailored for scene-specific language representation, requiring significant modeling time for each scene. This limits its applicability in tasks that demand rapid adaptation or broad generalization, such as in-the-wild scene understanding. Future work could focus on mitigating inherited biases and optimizing training pipelines to enhance scalability and generalization.

### v3

@@ 3.3 Super-Gaussian Clustering
$\widehat{\mathbf{e}}_{j}^{t+1}=\frac{1}{\sum_{i=1}^{N^{\prime}}\mathbb{I}(j\in\mathcal{N}_{i}){{\bm{A}}}_{ij}^{t}}\sum_{i=1}^{N^{\prime}}\mathbb{I}(j\in\mathcal{N}_{i}){{\bm{A}}}_{ij}^{t}\mathbf{e}_{i},$ (4)
@@
We optimize the SuperG clustering network to learn the association matrix $\bm{A}$}}, ensuring that the derived SuperG attributes $\widehat{\mathbf{e}}$ accurately reconstruct the anchor attributes $\mathbf{e}$. Note that $\mathbf{e}$ from Stage 1 (Section 3.2) are now frozen:
@@
$\mathcal{L}_{compact,\mathcal{X}}=\frac{1}{S}\sum_{j=1}^{S}\sum_{\mathbf{x}\in\mathcal{X}_{j}}\lVert\mathbf{x}-\widehat{{\bm{x}}}_{j}\rVert,$ (6)
@@ 4.1 Experimental Setup
Baselines and Metrics. We compare our method with representative NeRF-based and 3DGS-based baselines, including LERF [13], LangSplat [29], LEGaussian [34], and OpenGaussian [38]. For the open-vocabulary semantic segmentation task, CLIP-encoded text features are compared with rendered 2D language feature maps via cosine similarity to produce per-pixel semantic predictions [43], evaluated with mean Intersection over Union (mIoU) and mean Accuracy (mAcc). For the open-vocabulary object selection task, we perform text queries directly in 3D space [38], retrieving the most relevant SuperG s and rendering them into 2D for evaluation with mIoU and mAcc. Since NeRF is an implicit representation without explicit 3D positions, LERF cannot be applied to this task. We also report inference-time efficiency, measuring both runtime and memory consumption for text queries on trained 3D scenes. Specifically, we perform multiple queries from different viewpoints and report the average query time. We consider this metric particularly important for assessing the feasibility of deploying models on resource-constrained devices and enabling real-time querying in practical scenarios.
@@ 6 Acknowledgement
This work was supported by the German Research Foundation (DFG): SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP A1, project number: 276693517, and the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Siyun Liang.
@@
This supplementary document provides additional details about our method. Section A elaborates on the design of the Super-Gaussian and demonstrates its application to downstream tasks. Section B provides further implementation details, including the MLP architectures for neural Gaussian feature decoding and the adaptation of OpenGaussian for 2D open-vocabulary semantic segmentation comparison. Section C reports detailed efficiency analysis, including training time, inference speed, and memory consumption. Section D reports extended ablation studies on SuperG hyperparameters and module variants. Section E presents additional quantitative and qualitative results. Lastly, Section F discusses the limitations of our approach and outlines potential directions for future work.
@@ Appendix A Super-Gaussian Details
Figure 8: Example of Super-Gaussian graph. Each node represents a SuperG, connected to its $k$-nearest neighbors based on similarity in instance features. Through connected component analysis, the SuperG nodes are divided into two distinct instances. Within Instance 1, SuperG nodes are interconnected by similar hierarchical features, further splitting into two parts.
Module Design.
As shown in Figure 3, our SuperG clustering network consists of four learnable MLPs. Inspired by SPNet [10], we design three attribute-specific learnable functions $F_{\phi}$, $F_{\varphi}$, and $F_{\psi}$, each implemented as a single-hidden-layer MLP with ReLU activation. These functions independently encode the differences between an anchor and its $k$-nearest SuperG s in coordinates, segmentation features, and geometry, respectively, producing embeddings that reflect the relevance of each attribute for SuperG assignment. A final MLP, $F_{sg}$, then concatenates these embeddings and integrates spatial, semantic, and geometric cues into a probabilistic assignment.
@@
After training the SuperG association modules, we obtain a soft association map ${\bm{A}}\in\mathbb{R}^{N^{\prime}\times k}$. During inference, each anchor point is assigned to one of its $k$-nearest neighbors with the highest probability, leading to a hard SuperG assignment $\bar{{\bm{A}}}\in\mathbb{R}^{N^{\prime}\times 1}$. The attributes of each SuperG are then computed by averaging the attributes of its assigned anchors.
These SuperG s serve as the fundamental units for representing and interpreting the 3D scene. Specifically, as shown in Figure 8, we further construct a graph where nodes correspond to SuperG s. For instance segmentation, a node is connected to nodes within its $k$-nearest neighbors if their instance feature similarity exceeds a threshold $\tau_{Ins}$. Instances are then obtained via connected component analysis on this restricted graph. Similarly, part segmentation is achieved by building a SuperG graph within each instance and identifying connected components based on hierarchical feature similarity with threshold $\tau_{Hier}$. In practice, we set $k=3$, $\tau_{Ins}=0.8$, and $\tau_{Hier}=0.9$.
@@ Appendix B Additional Implementation Details
Figure 9: MLP structures for decoding different features.
@@
We employ MLPs to decode latent features, as shown in Figure 9. Each MLP contains a single hidden layer of dimension 32. Their decoding targets, however, differ: instance feature decoder $F_{I}$ and hierarchical feature decoder $F_{H}$ decode anchor-level features, while language feature decoder $F_{L}$ decodes SuperG-level features. Specifically, $F_{I}$ and $F_{H}$ take the anchor segmentation feature ${\bm{f}}^{s}$ and anchor position $\bm{x}$} as input, and predict the instance feature $\bm{g}$} and hierarchical feature $\bm{h}$} of the neural Gaussians spawned per anchor. In contrast, $F_{L}$ predicts the CLIP-aligned feature $\widehat{{\bm{l}}}$ for each SuperG, conditioned on its latent language feature $\widehat{{\bm{f}}}^{l}$ and its center $\widehat{{\bm{x}}}$.
@@
The contrastive instance feature loss $\mathcal{L}_{Ins}$, which is computed on the set of SAM-generated instance-level masks $\mathcal{M}$ and the rendered 2D instance feature map $\hat{{\bm{G}}}$, has been discussed in Section 3.2. Inspired by [41], we elaborate a similar yet more complicated contrastive hierarchical feature loss $\mathcal{L}_{Hier}$. This loss is defined on the part-level patches $\mathcal{P}=\{{{\bm{p}}}^{p}\in{\mathbb{R}}^{H\times W}\mid p=1,\dots,|\mathcal{P}|\}$ and the rendered 2D hierarchical feature map $\hat{{\bm{H}}}\in{\mathbb{R}}^{D_{h}\times H\times W}$.
@@
We use the SAM ViT-H model [16] to generate 2D masks from the input images and then extract language features for each instance mask using the OpenCLIP ViT-B/16 model following [29]. The training process is divided into three stages. In the first stage, we train the Scaffold-GS [23] with instance and hierarchical feature attributes for 30k iterations. In the second stage, we freeze the geometry and multi-granularity features network from stage one and train only the SuperG clustering network for another 30k iterations. Finally, in the last stage, we freeze all other parameters and optimize the language features for each SuperG for 10k iterations. The embedding dimensions for ${\bm{f}}^{g}$ and ${\bm{f}}^{s}$ are set to 32 [23], while instance and hierarchical features are 16-dimensional [41]. For optimization, we use the Adam [15] optimizer for the MLPs with an initial learning rate of 0.01 and an exponential annealing schedule of 0.001 as in [8].
@@
OpenGaussian [38] assigns language features to instance-level Gaussians, enabling direct language queries on 3D point clouds. However, this approach does not natively support 2D pixel-level semantic segmentation, making direct evaluation on ScanNet more challenging. To enable a fair comparison, we first identify category-relevant 3D Gaussians by iterating over all text prompts to predict language feature maps for open-vocabulary semantic segmentation. For each instance-level Gaussian cluster, we determine the corresponding text prompt ID and store these IDs in a label map, which is then used to generate the final semantic segmentation. By following this approach, occlusions at the instance-level Gaussians are not explicitly handled, leading to the occlusion artifacts observed in Figure 4.
@@
Table 4: Additional ablation studies on the LERF-OVS dataset [29] about the parameter choices for the SuperG Clustering Network. We use $s=1000$ SuperG s and $k=3$ for $k$-nearest neighbor in our implementation by default.
Appendix C Training and Inference Efficiency
@@
In Table 5, we report training time, inference time and memory consumption for LangSplat [29] and OpenGaussian [38] on the LERF-OVS dataset.
@@
When comparing our method to LangSplat and OpenGaussian across different training stages S1, S2, and S3, we find that our approach, while requiring longer training times in S1 and S2, achieves comparable efficiency in S3, indicating a trade-off due to the joint learning of instance and hierarchical features. GPU memory usage varies, with our method consuming more resources in S1 and S2 but significantly less in S3, showcasing the compact memory footprint enabled by our proposed SuperG s. Conversely, LangSplat exhibits consistent memory efficiency in S1 and S2, while OpenGaussian’s memory demands vary across different training stages. Most notably, our method consistently outperforms both baselines in inference speed across all scenarios, a critical advantage for real-time applications. This blend of rapid inference and flexible resource utilization highlights our method’s robustness for practical deployment in diverse computer vision tasks. In addition, our method supports multi-granularity scene understanding. Specifically, it enables semantic-level queries to retrieve groups of objects sharing the same language description, instance-level segmentation of a specific object, and further decomposition of this instance into fine-grained parts. In contrast, LangSplat [29] only supports semantic-level queries. OpenGaussian [38] extends to instance-level understanding but, similar to LangSplat, does not support finer-grained part segmentation. This demonstrates that our method provides a more comprehensive representation for 3D scene understanding.
@@ Appendix D Additional Ablation Studies
Super-Gaussian Clustering Approaches.
Given a pre-trained scene, our objective is to group anchors into meaningful SuperG s using their coordinates, segmentation features, and geometric properties. We explore two alternative approaches. First, we evaluate a simple $K$-means clustering algorithm by concatenating the aforementioned attributes and clustering them into $k=1000$ SuperG s. Second, we experiment with a traditional supervoxel generation approach. Specifically, we map the anchors to a point cloud, using the concatenated features as normals. We then apply the Voxel Cloud Connectivity Segmentation (VCCS) algorithm [26] to compute the SuperG s. Finally, we compare these two non-learning-based approaches with our learning-based method for Anchor-to-SuperG association.
@@
Table 6: Ablation study of SuperG clustering approaches on the teatime scene of LERF-OVS.
@@
Table 7: Ablation study on the SuperG clustering network and its components on the teatime scene of LERF-OVS.
@@
As introduced in Section A, we employ three MLPs $F_{\phi}$, $F_{\varphi}$, and $F_{\psi}$ to capture the coordinate, segmentation, and geometric relationships between anchors and their $k$-nearest neighbors. To evaluate the contributions of these MLPs, we conduct an ablation study. Notably, in experiments where none of these MLPs are used, we directly concatenate the differences of the attributes as input to $F_{sg}$ for predicting the association matrix. The results presented in Table 7 demonstrate that each MLP contributes to improving the SuperG assignments. The MLP $F_{\varphi}$, which accounts for the segmentation feature differences between the anchor and the SuperG, has the most significant impact. In particular, using only $F_{\varphi}$ yields a relatively high mIoU, emphasizing its effectiveness in aligning semantic features. However, our full setup that integrates the $F_{\phi}$ and $F_{\psi}$ for coordinate and geometric feature information further enhances mAcc. This suggests that incorporating additional spatial and geometric context refines the SuperG assignments, leading to a more precise understanding of the scene.
@@ Appendix E Additional Results
Qualitative Results in Occlusion Cases. As illustrated in Figure 10, our method queries objects directly in 3D space, effectively mitigating occlusion issues (e.g., the bear leg under the table can be retrieved). Moreover, the queried objects exhibit multi-view consistency, enabling comprehensive scene understanding in 3D. For additional qualitative results, please refer to the accompanying videos.
Additional Quantitative Results on ScanNet.
We report the results on more categories in the ScanNet dataset in Table 8.
@@ Appendix F Limitations
Despite the advancements achieved by our method, certain limitations remain. First, our approach inherits biases from the original visual foundation models, which may constrain performance and limit generalization to diverse or unseen scenarios. Second, our method is tailored for scene-specific language representation, requiring significant modeling time for each scene. This limits its applicability in tasks that demand rapid adaptation or broad generalization, such as in-the-wild scene understanding. Future work could focus on mitigating inherited biases and optimizing training pipelines to enhance scalability and generalization.